OpenStudy (anonymous):

Let F(X)= x/2 and g(x)=2f(x) . Then g(2) = _____.

OpenStudy (anonymous):

\[g(2)=2\times f(2)\]

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

we are not done. do you know what \(f(2)\) is?

OpenStudy (anonymous):

Eh no :( im horrible at math :/

OpenStudy (anonymous):

itis function

OpenStudy (anonymous):

\(f(x)=\frac{2}{x}\) means take your number and divide it by 2 so \(f(2)=\frac{2}{2}=1\)

OpenStudy (anonymous):

and therefore \(g(2)=2\times f(2)=2\times 1=2\)

OpenStudy (anonymous):

OOOOOOO I get 2 over the number

OpenStudy (anonymous):

well, not two over the number, but rather the number over (divided by ) 2

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

yw